# Assessing Adaptive World Models in Machines with Novel Games

### v1

@@ Abstract
Human intelligence exhibits a remarkable capacity for rapid adaptation and effective problem-solving in novel and unfamiliar contexts. We argue that this profound adaptability is fundamentally linked to the efficient construction and refinement of internal representations of the environment, commonly referred to as world models, and we refer to this adaptation mechanism as world model induction. However, current understanding and evaluation of world models in artificial intelligence (AI) remains narrow, often focusing on static representations learned from training on a massive corpora of data, instead of the efficiency and efficacy of models in learning these representations through interaction and exploration within a novel environment. In this Perspective, we provide a view of world model induction drawing on decades of research in cognitive science on how humans learn and adapt so efficiently; we then call for a new evaluation framework for assessing adaptive world models in AI. Concretely, we propose a new benchmarking paradigm based on suites of carefully designed games with genuine, deep and continually refreshing novelty in the underlying game structures — we refer to this kind of games as novel games. We detail key desiderata for constructing these games and propose appropriate metrics to explicitly challenge and evaluate the agent’s ability for rapid world model induction. We hope that this new evaluation framework will inspire future evaluation efforts on world models in AI and provide a crucial step towards developing AI systems capable of the human-like rapid adaptation and robust generalization — a critical component of artificial general intelligence.
@@ 1 Introduction
To drive AI progress towards this goal, and to be able to measure that progress, we argue for a comprehensive evaluation framework grounded in the cognitive science theories and experimental paradigms that have been used to study world model induction in humans. We propose that games are a uniquely advantageous domain for evaluating these capabilities in AI, given their inherently rich, often hierarchical structures in concepts and skills and their well-controlled environments. Concretely, we introduce an evaluation paradigm centered around the concept of novel games. Within this framework, a novel games is defined not simply by unseen instances or parametric variations within a familiar game structure, but by environments with structured novelty, where underlying rules, mechanics, object properties, or objectives are initially unknown, hidden, or dynamically changing. Success requires agents to rapidly infer these latent dynamics and causal structures through active, limited interaction and exploration, effectively performing world model induction on-the-fly.
@@ 2 World Models in Humans and Machines
A precise definition of world models is often debated (Ding et al.,, 2024). For the purposes of this paper, we define world model as an agent’s internal representation of an environment, including its dynamics, rules, objects, and underlying causal relations. The utility of such a model lies in its ability to allow agents to efficiently simulate different world states for effective decision making, planning, and problem solving. For instance, empirically, compared to model-free reinforcement learning, model-based reinforcement learning has shown to be more data-efficient and affords better generalization (Kaiser et al.,, 2019; Moerland et al.,, 2023).
@@
How do humans achieve such remarkable sample efficiency in learning about the world? Extensive research in cognitive science have formulated human learning and intuitive theory induction within a hierarchical Bayesian framework (Gopnik and Wellman,, 2012; Gelman and Legare,, 2011; Tenenbaum et al.,, 2011; Ullman and Tenenbaum,, 2020), where theories and concepts are learned and represented at different levels of abstraction. We can draw a similar conceptual framework for world models, as shown in Figure 1.
@@
Learning can be understood as construction and refinement of such hierarchical world model. When an agent receives new observations ($e$), its beliefs about the underlying world model $\Omega$ are updated. The posterior probability distribution over possible world models, given new data $e$, is computed by inverting the generative model:
@@
The efficacy and sample efficiency of this Bayesian update process are significantly enhanced when the agent is not merely a passive observer but actively seeks out informative data. Insights from cognitive science and developmental psychology, particularly the “the child as scientist” framework (Schulz,, 2012; Gopnik and Wellman,, 1992; Gopnik,, 1996), suggest that human learning is characterized by hypothesis-driven exploration. Starting at infancy, humans learning involves actively planning and designing “experiments” — actions that (intentionally or not) effectively generate informative observations that can discriminate between competing hypotheses about the underlying world model $\Omega$. We expect that this kind of active hypothesis-driven approach to generating data is critical for any intelligent system to converge on an accurate representation of a new, unfamiliar domain with minimal interaction and observation, and more generally for sample-efficient world model induction.
@@
    Cross-Domain Generalization and Meta-Learning: The development of meta-learning capabilities, enabling faster and more efficient adaptation to new domains by leveraging prior world models. This reflects the human ability to build generalizable knowledge (e.g., intuitive physics, intuitive psychology) that can bootstrap learning and generalize broadly to new tasks, even those with fundamentally different domain characteristics (Spelke and Kinzler,, 2007; Allen et al.,, 2020; Lake et al.,, 2017; Chollet,, 2019). However, one also needs to decide what aspects of previous world models are relevant for constructing new world models for a new domain. For example, humans have a general intuitive theory of physics that can broadly be applied to reasoning about physics of objects in any new domain, but humans can also very effectively learn and interact with novel environments that violate basic physical properties (Liu and Xu,, 2024).
Rapid learning and robust generalization function as complementary components contributing to overall adaptive intelligence and map conceptually onto different facets of the hierarchical Bayesian world model induction formulation. Rapid learning in a new domains focuses on the efficiency of the inference process itself – how quickly the agent can leverage new observations $e$ to compute and refine the posterior distribution over world models $P(\Omega|e)$ and converge on an accurate world model $\Omega_{1}$ for a new environment. Robust generalization contributes to a more informative prior $P(\Omega_{1}|\Omega_{2},...\Omega_{n})$ over plausible world models, biasing learning towards likely structures for the new task based on prior experiences and learned abstract principles, from the same (within-domain) or other domains (cross-domain).
@@
These limitations highlight a critical need for a new evaluation framework designed to comprehensively assess world model in machines. In this perspective, we call for future AI evaluation efforts to holistically assess world models in machines according to the framework outlined above. We contend that games provide particularly rich and controlled environments uniquely well-suited for systematically evaluating rapid model adaptation and the process of world model induction. The inherent structure of many games, with concepts, rules, and skills often organized hierarchically from low-level actions to high-level strategies, aligns well with the requirement for evaluating the learning of world models at different levels of abstraction (See Fig. 1). The remainder of this paper details how games can serve this purpose and proposes a game-based benchmark proposal in subsequent sections.
@@ 4.1.1 Desideratum 1: Novelty in Game Structures
We encourage designing game environments where the underlying mechanics are not fully transparent or pre-specified to the player. Instead, crucial aspects – including types of objects, specific rules governing interactions, affordances and properties of objects, or the consequences of actions – should be partially or entirely latent, requiring the agent to infer them through active gameplay, exploration, and experimentation. We highlight three such game environments from previous studies in Figure 2.
@@ 4.1.3 Desideratum 3: Diversity in World Models and Learning Mechanisms
To span the diverse array of challenges people—and AI systems in a human-world may face—the benchmark game suite should encompass significant diversity to necessitate different types of world models agents are compelled to induce. For example, while some games may primarily involve learning about spatial relationships or object physics, others can require understanding and modeling other agents in multi-agent games (whether competitive or collaborative with other human or artificial agents). Successfully adapting in multi-agent scenarios often requires agents to develop sophisticated mental models about other agents, inferring and representing their goals, beliefs, intentions, or emotional states (often referred to as Theory of Mind (Gopnik and Wellman,, 1992)), which constitutes a crucial aspect of human social adaptation.
@@ 4.2 Evaluation of AI Agent’s World Modeling Capacity
We feature the work by Tsividis et al., (2021) in Figure 3a as a case study. In this example, they compute learning efficiency scores of different models and plot the success rate of the models at all game levels as a function of learning samples, which allows direct comparison of sample efficiency of humans vs models at learning and solving new tasks.
@@
Beyond quantitative metrics, a qualitative analysis of how agents explore and learn within novel environments can reveal crucial insights into their capabilities for world model induction. Different learning approaches often manifest in distinct exploration patterns, reflecting their strategies for gathering information and inferring rules. As demonstrated in Fig 3b from Tsividis et al., (2021), human players tend to exhibit targeted and efficient exploration, focusing on areas relevant to understanding the game mechanics. Similarly, the EMPA model proposed in the same paper shows more structured but sometimes less efficient exploration, while DDQN (Van Hasselt et al.,, 2016) often displays highly diffuse and less directed exploration, indicative of a struggle to efficiently form coherent internal representations of the environment. Such qualitative visual comparisons can highlight the differing capacities of agents to induce and utilize effective world models, revealing whether they are genuinely understanding the game’s rules or merely optimizing for short-term rewards.
@@
To properly assess rapid world model induction, it is essential to gain insight into the nature of the internal representations the agent constructs and how these representations are dynamically updated in response to new observations and actions.
@@
For neural networks, inspecting the internal world model is more complex, typically involving analyzing representation spaces, activation patterns, attention mechanisms, or tracing reasoning processes through the network (Vafa et al.,, 2024). Techniques such as probing specific network layers for learned features related to game mechanics or dynamics can reveal aspects of the implicit world model. For many of the large foundation models that can interact with humans with natural language, we can also examine their understanding of the game mechanics through targeted question answering at different levels of abstraction. For instance, one may ask questions about abstract principles, such as knowledge about objects (e.g. object permanence) or properties specific to the game, such as the mass of a particular object.
@@ 5.1 Game Design and Creation
One approach is to design every novel game entirely from scratch, which, while offering maximum control over the novelty injected, can be a significantly labor-intensive process, potentially limiting the scale and diversity of the benchmark suite. For instance,  Zhang et al., (2024) manually created $120$ variants of grid-based games, like Tic-Tac-Toe and Gomoku. More broadly, drawing inspiration from human creativity, which often builds upon existing structures to create novel variants (Vélez et al.,, 2024; Zhao et al.,, 2024; Youn et al.,, 2015; Collins et al.,, 2025), AI benchmark creators can adopt a similar strategy. Instead of solely focusing on generating games disconnected from known forms, a promising avenue involves creating interesting and challenging novel variants by systematically altering the rules, objectives, player counts, or interaction dynamics of existing popular games. This method leverages the existing human familiarity with the base game, providing a potentially more intuitive starting point for both human and AI adaptation to the introduced changes (we highlight some examples of systematic alteration in Fig. 4).
@@ 5.2 Human Evaluation and Validation
Human input is essential for ensuring that the novel games generated for the benchmark meet the desideratum we lay out above: of being intuitive, adaptable, and learnable for people. Benchmark creators can employ a combination of self-reported data and behavioral measures as proxies for these qualities. For instance, soliciting ratings from human players regarding game engagement, playability, and perceived difficulty can provide valuable qualitative feedback. Additionally, analyzing the playing traces and learning curves of human participants can offer quantitative insights into their ability to infer rules, develop strategies, and adapt to the novel mechanics, thereby confirming that the games are indeed human-learnable and intuitive as intended. For instance, ARC-AGI conducts human evaluations to make sure benchmark tasks can be solved by most human players (Chollet et al.,, 2025). In addition to controlled human experiments, one could envision the creation of dedicated game platforms where human players can engage at their own time or where human players can voluntarily interact with or compete against AI models, similar to existing initiatives like TextArena (Guertler et al.,, 2025). This approach offers the advantage of collecting large-scale data in a relatively naturalistic setting. Such human data can be useful for benchmarking AI on these novel games and also allow researchers to perform detailed analysis to investigate and compare the process of world model induction in AI systems versus humans.
@@ 6 Discussion and Looking Forward
While we believe this paradigm offers a valuable path forward, we acknowledge several important outstanding questions and challenges that warrant future investigation and refinement. A fundamental question that may arise regarding our central thesis is the extent to which hierarchical and adaptive world models are truly necessary for rapid and efficient adaptation. Our Perspective is strongly motivated by and aligns with extensive evidence from cognitive science and developmental psychology (Tenenbaum et al.,, 2011; Gopnik and Wellman,, 2012), which highlights the crucial role of constructing and refining internal models in human learning and adaptation. However, one may argue that alternative computational mechanisms could facilitate adaptation without relying on world models at all (Brooks,, 1991). For example, one may question if a reactive or purely end-to-end system with massive capacity might exhibit surprising levels of adaptation through pattern generalization across diverse training data, even if they lack a structured, predictive model of environmental dynamics. This view has met some skepticism by recent work in AI, which argues that world models are essential for building general agents (Richens et al.,, 2025). Nevertheless, the extent to which such model-free approaches can achieve human-level rapid, sample-efficient adaptation and robust generalization remains an open empirical question.
@@
Lastly, there are important open questions about the external validity of findings from novel game benchmarks. While games provide a valuable, controlled, and complex environment for studying aspects of adaptive intelligence and world model induction, it remains an open question how directly the adaptive capabilities demonstrated within these simulated environments will transfer to the complexities, continuous nature, and richness of real-world adaptation challenges. Rigorously assessing and quantifying the external validity of benchmark results — for instance, by designing complementary evaluations in limited real-world scenarios or high-fidelity, open-ended simulation environments grounded in physical or social realism — is a critical direction for future research. Such work is necessary to ensure that progress on game benchmarks translates meaningfully to progress on real-world AI capabilities, particularly if AI systems engage with people (Collins et al.,, 2024).

### v2

@@ Abstract
Human intelligence exhibits a remarkable capacity for rapid adaptation and effective problem-solving in novel and unfamiliar contexts. We argue that this profound adaptability is fundamentally linked to the efficient construction and refinement of internal representations of the environment, commonly referred to as world models, and we refer to this adaptation mechanism as world model induction. However, current understanding and evaluation of world models in artificial intelligence (AI) remains narrow, often focusing on static representations learned from training on massive corpora of data, instead of the efficiency and efficacy in learning these representations through interaction and exploration within a novel environment. In this Perspective, we provide a view of world model induction drawing on decades of research in cognitive science on how humans learn and adapt so efficiently; we then call for a new evaluation framework for assessing adaptive world models in AI. Concretely, we propose a new benchmarking paradigm based on suites of carefully designed games with genuine, deep and continually refreshing novelty in the underlying game structures — we refer to this class of games as novel games. We detail key desiderata for constructing these games and propose appropriate metrics to explicitly challenge and evaluate the agent’s ability for rapid world model induction. We hope that this new evaluation framework will inspire future evaluation efforts on world models in AI and provide a crucial step towards developing AI systems capable of human-like rapid adaptation and robust generalization — a critical component of artificial general intelligence.
@@ 1 Introduction
To drive AI progress towards this goal, and to be able to measure that progress, we argue for a comprehensive evaluation framework grounded in the cognitive science theories and experimental paradigms that have been used to study world model induction in humans. We propose that games are a uniquely advantageous domain for evaluating these capabilities in AI, given their inherently rich, often hierarchical structures in concepts and skills and their well-controlled environments. Concretely, we introduce an evaluation paradigm centered around the concept of novel games. Within this framework, a novel game is defined not simply by unseen instances or parametric variations within a familiar game structure, but by environments with structured novelty, where underlying rules, mechanics, object properties, or objectives are initially unknown, hidden, or dynamically changing. Success requires agents to rapidly infer these latent dynamics and causal structures through active, limited interaction and exploration, effectively performing world model induction on-the-fly.
@@ 2 World Models in Humans and Machines
The precise definition of world models is often debated (Ding et al.,, 2024). For the purposes of this paper, we define world model as an agent’s internal representation of an environment, including its dynamics, rules, objects, and underlying causal relations. The utility of such a model lies in its ability to allow agents to efficiently simulate different world states for effective decision making, planning, and problem solving. For instance, empirically, compared to model-free reinforcement learning, model-based reinforcement learning has shown to be more data-efficient and affords better generalization (Kaiser et al.,, 2019; Moerland et al.,, 2023).
@@
How do humans achieve such remarkable sample efficiency in learning about the world? Extensive research in cognitive science has formulated human learning and intuitive theory induction within a hierarchical Bayesian framework (Gopnik and Wellman,, 2012; Gelman and Legare,, 2011; Tenenbaum et al.,, 2011; Ullman and Tenenbaum,, 2020), where theories and concepts are learned and represented at different levels of abstraction. We can draw a similar conceptual framework for world models, as shown in Figure 1.
@@
Learning can be understood as the construction and refinement of such hierarchical world model. When an agent receives new observations ($e$), its beliefs about the underlying world model $\Omega$ are updated. The posterior probability distribution over possible world models, given new data $e$, is computed by inverting the generative model:
@@
The efficacy and sample efficiency of this Bayesian update process are significantly enhanced when the agent is not merely a passive observer but actively seeks out informative data. Insights from cognitive science and developmental psychology, particularly the “the child as scientist” framework (Schulz,, 2012; Gopnik and Wellman,, 1992; Gopnik,, 1996), suggest that human learning is characterized by hypothesis-driven exploration. Starting at infancy, human learning involves actively planning and designing “experiments” — actions that (intentionally or not) effectively generate informative observations that can discriminate between competing hypotheses about the underlying world model $\Omega$. We expect that this kind of active hypothesis-driven approach to generating data is critical for any intelligent system to converge on an accurate representation of a new, unfamiliar domain with minimal interaction and observation, and more generally for sample-efficient world model induction.
@@
    Cross-Domain Generalization and Meta-Learning: The development of meta-learning capabilities, enabling faster and more efficient adaptation to new domains by leveraging prior world models. This reflects the human ability to build generalizable knowledge (e.g., intuitive physics, intuitive psychology) that can bootstrap learning and generalize broadly to new tasks, even those with fundamentally different domain characteristics (Spelke and Kinzler,, 2007; Allen et al.,, 2020; Lake et al.,, 2017; Chollet,, 2019). However, one also needs to decide which aspects of previous world models are relevant for constructing new world models for a new domain. For example, humans have a general intuitive theory of physics that can broadly be applied to reasoning about physics of objects in any new domain, but humans can also very effectively learn and interact with novel environments that violate basic physical properties (Liu and Xu,, 2024).
Rapid learning and robust generalization function as complementary components contributing to overall adaptive intelligence and map conceptually onto different facets of the hierarchical Bayesian world model induction formulation. Rapid learning in a new domain focuses on the efficiency of the inference process itself – how quickly the agent can leverage new observations $e$ to compute and refine the posterior distribution over world models $P(\Omega|e)$ and converge on an accurate world model $\Omega_{1}$ for a new environment. Robust generalization contributes to a more informative prior $P(\Omega_{1}|\Omega_{2},...\Omega_{n})$ over plausible world models, biasing learning towards likely structures for the new task based on prior experiences and learned abstract principles, from the same (within-domain) or other domains (cross-domain).
@@
These limitations highlight a critical need for a new evaluation framework designed to comprehensively assess world model in machines. In this perspective, we call for future AI evaluation efforts to holistically assess world models in machines according to the framework outlined above. We contend that games provide particularly rich and controlled environments uniquely well-suited for systematically evaluating rapid model adaptation and the process of world model induction. The inherent structure of many games, with concepts, rules, and skills often organized hierarchically from low-level actions to high-level strategies, aligns well with the requirement for evaluating the learning of world models at different levels of abstraction (See Fig. 1). The remainder of this paper details how games can serve this purpose and proposes a new game-based benchmark proposal in subsequent sections.
@@ 4.1.1 Desideratum 1: Novelty in Game Structures
We encourage the design of game environments where the underlying mechanics are not fully transparent or pre-specified to the player. Instead, crucial aspects – including types of objects, specific rules governing interactions, affordances and properties of objects, or the consequences of actions – should be partially or entirely latent, requiring the agent to infer them through active gameplay, exploration, and experimentation. We highlight three such game environments from previous studies in Figure 2.
@@ 4.1.3 Desideratum 3: Diversity in World Models and Learning Mechanisms
To span the diverse array of challenges people—and AI systems in a human-world may face—the benchmark game suite should encompass significant diversity to necessitate different types of world models that agents are compelled to induce. For example, while some games may primarily involve learning about spatial relationships or object physics, others can require understanding and modeling other agents in multi-agent games (whether competitive or collaborative with other human or artificial agents). Successfully adapting in multi-agent scenarios often requires agents to develop sophisticated mental models about other agents, inferring and representing their goals, beliefs, intentions, or emotional states (often referred to as Theory of Mind (Gopnik and Wellman,, 1992)), which constitutes a crucial aspect of human social adaptation.
@@ 4.2 Evaluation of AI Agent’s World Modeling Capacity
We feature the work by Tsividis et al., (2021) in Figure 3a as a case study. In this example, they compute learning efficiency scores of different models and plot the success rate of the models at all game levels as a function of learning samples, which allows direct comparison between the sample efficiency of humans vs models at learning and solving new tasks.
@@
Beyond quantitative metrics, a qualitative analysis of how agents explore and learn within novel environments can reveal crucial insights into their capabilities for world model induction. Different learning approaches often manifest in distinct exploration patterns, reflecting their strategies for gathering information and inferring rules. As demonstrated in Figure 3b from Tsividis et al., (2021), human players tend to exhibit targeted and efficient exploration, focusing on areas relevant to understanding the game mechanics. Similarly, the EMPA model proposed in the same paper shows more structured but sometimes less efficient exploration, while DDQN (Van Hasselt et al.,, 2016) often displays highly diffuse and less directed exploration, indicative of a struggle to efficiently form coherent internal representations of the environment. Such qualitative visual comparisons can highlight the differing capacities of agents to induce and utilize effective world models, revealing whether they are genuinely understanding the game’s rules or merely optimizing for short-term rewards.
@@
To properly assess rapid world model induction, it is essential to gain insight into the nature of the internal representations that the agent constructs and how these representations are dynamically updated in response to new observations and actions.
@@
For neural networks, inspecting the internal world model is more complex, typically involving analyzing representation spaces, activation patterns, attention mechanisms, or tracing reasoning processes through the network (Vafa et al.,, 2024). Techniques such as probing specific network layers for learned features related to game mechanics or dynamics can reveal aspects of the implicit world model. For many of the large foundation models that can interact with humans in natural language, we can also examine their understanding of the game mechanics through targeted question answering at different levels of abstraction. For instance, one may ask questions about abstract principles, such as knowledge about objects (e.g. object permanence) or properties specific to the game, such as the mass of a particular object.
@@ 5.1 Game Design and Creation
One approach is to design every novel game entirely from scratch, which, while offering maximum control over the novelty injected, can be a significantly labor-intensive process, potentially limiting the scale and diversity of the benchmark suite. For instance,  Zhang et al., (2024) manually created $120$ variants of grid-based games, such as Tic-Tac-Toe and Gomoku. More broadly, drawing inspiration from human creativity, which often builds upon existing structures to create novel variants (Vélez et al.,, 2024; Zhao et al.,, 2024; Youn et al.,, 2015; Collins et al.,, 2025), AI benchmark creators can adopt a similar strategy. Instead of solely focusing on generating games disconnected from known forms, a promising avenue involves creating interesting and challenging novel variants by systematically altering the rules, objectives, player counts, or interaction dynamics of existing popular games. This method leverages the existing human familiarity with the base game, providing a potentially more intuitive starting point for both human and AI adaptation to the introduced changes (we highlight some examples of systematic alteration in Fig. 4).
@@ 5.2 Human Evaluation and Validation
Human input is essential for ensuring that the novel games generated for the benchmark meet the desiderata we lay out above: of being intuitive, adaptable, and learnable for people. Benchmark creators can employ a combination of self-reported data and behavioral measures as proxies for these qualities. For instance, soliciting ratings from human players regarding game engagement, playability, and perceived difficulty can provide valuable qualitative feedback. Additionally, analyzing the playing traces and learning curves of human participants can offer quantitative insights into their ability to infer rules, develop strategies, and adapt to the novel mechanics, thereby confirming that the games are indeed human-learnable and intuitive as intended. For instance, ARC-AGI conducts human evaluations to make sure benchmark tasks can be solved by most human players (Chollet et al.,, 2025). In addition to controlled human experiments, one could envision the creation of dedicated game platforms where human players can engage at their convenience or where human players can interact with or compete against AI models, similar to existing initiatives like TextArena (Guertler et al.,, 2025). This approach offers the advantage of collecting large-scale data in a relatively naturalistic setting. Such human data can be useful for benchmarking AI on these novel games and also allow researchers to perform detailed analyses to investigate and compare the process of world model induction in AI systems versus humans.
@@ 6 Discussion and Looking Forward
While we believe this paradigm offers a valuable path forward, we acknowledge several important outstanding questions and challenges that warrant future investigation and refinement. A fundamental question that may arise regarding our central thesis is the extent to which hierarchical and adaptive world models are truly necessary for rapid and efficient adaptation. Our Perspective is strongly motivated by and aligned with extensive evidence from cognitive science and developmental psychology (Tenenbaum et al.,, 2011; Gopnik and Wellman,, 2012), which highlights the crucial role of constructing and refining internal models in human learning and adaptation. However, one may argue that alternative computational mechanisms could facilitate adaptation without relying on world models at all (Brooks,, 1991). For example, one may question if a reactive or purely end-to-end system with massive capacity might exhibit surprising levels of adaptation through pattern generalization across diverse training data, even if it lacks a structured, predictive model of environmental dynamics. This view has met with some skepticism by recent work in AI, which argues that world models are essential for building general agents (Richens et al.,, 2025). Nevertheless, the extent to which such model-free approaches can achieve human-level rapid, sample-efficient adaptation and robust generalization remains an open empirical question.
@@
Lastly, there are important open questions about the external validity of findings from novel game benchmarks. While games provide a valuable, controlled, and complex environment for studying aspects of adaptive intelligence and world model induction, it remains an open question how directly the adaptive capabilities demonstrated within these simulated environments will transfer to the complexities, continuous nature, and richness of real-world adaptation challenges. Rigorously assessing and quantifying the external validity of benchmark results — for instance, by designing complementary evaluations in limited real-world scenarios or high-fidelity, open-ended simulation environments grounded in physical or social realism — is a critical direction for future research. Such work is necessary to ensure that progress on game benchmarks translates meaningfully to progress on real-world AI capabilities, particularly if AI systems are designed to engage with people (Collins et al.,, 2024).